Grade 128 grade maths

Factorise the expressions and divide them as directed.
(m^2 –14m–32)÷(m+2)

Pawan Prajapati
5 Years agoGrade 12

1 Answer

Harshit Singh
5 Years ago
Dear Student

Solve for m^2–14m–32, we have
m^2–14m–32 = m^2+2m-16m–32
= m(m+2)–16(m+2) = (m–16)(m+2)
Now, (m^2–14m–32)÷(m+2) = (m–16)(m+2)/(m+2) = m-16

Thanks